A movie theater charges $ 8 for adults and $ 6 for seniors. On a particular day when 344 people paid an admission, the total receipts were $ 2575. Round your answers to the nearest integers.
Answer by Boreal(15235)

adults=x
seniors=y
x+y=344
8x+6y=2575
-6x-6y= -2064, multiplying the top equation by (-6)
2x=511
x=255.5 or 256 adults
y=88 seniors
256*$8=$2048
88*$6=$528
$2576
===